# Boaz Moerman

### v2

@@ 6. Quasi-Campana points and the log-canonical divisor
In this section we specialize the Conjecture 1.4 to Campana points, weak Campana points and Darmon points as defined in Definition 3.23. In particular, we will clarify the relation of Conjecture 1.4 with the conjecture [PSTVA21, Conjecture 1.1] on Campana points of bounded height. In order to uniformly discuss these different kinds of points, we introduce the notion of a quasi-Campana pair.
@@
Let $(X,M)$ be a smooth proper pair over a field $K$. Assume that that there exists a log-canonical class $K_{(X,M),\log}$ for $(X,M)$. Let $L$ be a big and nef $\mathbb{Q}$-divisor class on $X$. Then
@@
The assumption on the effective cone in the Proposition is satisfied for many varieties $X$, including Mori dream spaces such as Fano varieties [BCHM10, Corollary 1.3.2].
@@
This conjecture is a related to a conjecture by Campana [Cam11, Conjecture 9.10] on Campana rational connectedness for Campana pairs (see also [CLT24, Conjecture 1.4]), which implies Conjecture 6.13 for any pair $(X,M)$ corresponding to the Campana points on $(X,D_{\mathbf{m}})$.
@@ 7.1. The fundamental group
We will introduce the fundamental group of a smooth pair using the valuations $v_{(\mathbf{m},c)}$ introduced in Definition 4.12. For a smooth variety $X$ over a field $K$, the étale fundamental group $\pi_{1}(X)$ admits a description using the valuations of function field. Indeed, if $f\colon Y\rightarrow X$ is a finite morphism of smooth varieties, then purity of the ramification locus [Stacks, Tag 0BMB] implies that $f$ is étale if and only if it is unramified at every prime divisor $D$ of $X$.
@@
If $(X,M)$ is a smooth pair corresponding to Darmon points, then [Moe24, Lemma 8.7.] and purity of the branch locus [Stacks, Tag 0BMB] together imply that the fundamental group $\pi_{1}(X,M)$ is the étale fundamental group of the corresponding root stack.
@@ 7.2. The Brauer group of a pair
In [MiNaSt22], Mitankin, Nakahara and Streeter introduced the Brauer group $\operatorname{Br}(X,D_{\mathbf{m}})$ for Campana pairs $(X,D_{\mathbf{m}})$ in their study of Darmon points. This Brauer group was then used by Chow, Loughran, Takloo-Bighash and Tanimoto in their prediction [CLTT24, Conjecture 8.3] for the leading constant in the analogue of Manin’s conjecture for Campana points. This group is defined as
@@
where $\partial_{D_{i}}$ is the Witt residue as defined in [CoSk21, Definition 1.4.11(ii)]. Note that this is the Brauer group of the corresponding root stack [MiNaSt22, Remark 3.15].
@@
is $\partial_{\mathbf{m},c}=\gcd(\mathbf{m})\cdot\partial_{\mathbf{m}/\gcd(\mathbf{m}),c}$, where $K_{(\mathbf{m}/\gcd(\mathbf{m}),c)}$ is the residue field of the valuation $v_{(\mathbf{m}/\gcd(\mathbf{m}),c)}$ and $\partial_{\mathbf{m}/\gcd(\mathbf{m}),c}$ is the residue corresponding to the valuation $v_{\mathbf{m}/\gcd(\mathbf{m}),c}$, as defined in [CoSk21, Definition 1.4.11(ii)].
@@ 7.3. The algebraic Brauer group for pairs over toric varieties
In this section we describe the algebraic Brauer group for pairs over toric varieties in terms of automorphic characters on the open torus, which we will use in Section 9.5 and in [Moe25toric].
Let $(X,M)$ be a smooth pair such that $X$ is a toric variety over a number field and such that the divisors $D_{1},\dots,D_{n}$ are the torus-invariant prime divisors, defined over the splitting field $E/K$ of the dense torus. Note that such a pair is always divisorial, as any nonempty intersection of torus-invariant prime divisors is a toric variety itself and thus irreducible. By using the correspondence between Brauer group elements on a torus and automorphic characters as in [Lou18], we will describe the algebraic Brauer group $\operatorname{Br}_{1}(X,M)$ in terms of characters if $X$ is a (rational) toric variety and the divisors are the torus invariant divisors.
@@
In particular, given an automorphic character $\chi\colon U(\mathbf{A}_{K})\rightarrow S^{1}$, we obtain an automorphic character $\chi_{\mathbf{m}}\colon T_{\mathbf{m}}(\mathbf{A}_{K})\rightarrow S^{1}$ for every $\mathbf{m}\in\mathbb{N}_{\textnormal{red}}/G$. The following lemma generalizes [ShSt24, Lemma 3.25].
@@
Here $\B(U)=\ker(\operatorname{Br}_{1}(U)\rightarrow\prod_{v\in\Omega_{K}}\operatorname{Br}_{1}(U_{K_{v}}))$, which is trivial when $X$ is rational. This isomorphism identifies the Brauer-Manin pairing with the pairing between $U(\mathbf{A}_{K})$ and the corresponding automorphic characters.
@@
For a torus $T$ over $K$ we write $\operatorname{Br}_{e}T=\{b\in\operatorname{Br}_{1}U\mid 1_{T}^{*}b=0\}$, where $1_{T}\colon\operatorname{Spec}K\rightarrow T$ is the identity element, so that we have $\operatorname{Br}_{1}T\cong\operatorname{Br}K\times\operatorname{Br}_{e}T$. We similarly write $\operatorname{Br}_{e}(X,M)=\operatorname{Br}(X,M)\cap\operatorname{Br}_{1}U$.
@@
These observations combine with [Lou18, Lemma 4.7] to give a commutative diagram
@@
Finally, the triviality of $\B(U)$ when $X$ is rational is proven in [Lou18, Corollary 4.6]. ∎
@@ 8. The leading constant
In this section we give a description for the leading constant in Conjecture 1.4 when the divisor class $L$ is adjoint rigid and the fundamental group $\pi(X_{\overline{K}},M)$ of the pair is abelian. Our expression for the leading constant will involve the Fujita-invariant and the $b$-invariant, as well as analogues of the $\alpha$-constant and the Tamagawa constant for pairs. In Manin’s conjecture, the complement of the support of the adjoint divisor is used to define several of these invariants.
The $\alpha$-constant and the Tamagawa constant in Manin’s conjecture, as formulated in [BaTs98], are defined using the complement $X^{\circ}$ of the adjoint divisor. Our conjecture replaces the role of this open subvariety with a pair $(X^{\circ},M^{\circ})\subset(X,M)$.
@@
where $\tau_{X,v}$ is the measure on $X(K_{v})$ as defined in [CLTs10, §2.1.8] induced by the chosen metrisation of $K_{X}$. Note that the measure $\tau_{(X^{\circ},M^{\circ})}$ can be viewed as an analogue of the Tamagawa measure defined in [CLTs10, Theorem 1.1].
As in [CLTT24, Conjecture 8.3], we define the Tamagawa constant as a sum over Brauer classes in $\operatorname{Br}(X^{\circ},M^{\circ})/\operatorname{Br}K$.
@@
Similarly to the constant in [CLTT24, Conjecture 8.3], it is not clear whether the sum defining the Tamagawa constant converges, and if it does, whether it converges absolutely. In the case that the sum does not converge absolutely, the sum $\sum_{b\in\operatorname{Br}(X^{\circ},M^{\circ})/\operatorname{Br}K}\hat{\tau}(b)$ should be interpreted as the limit
@@
where $\operatorname{ev}_{b}(x)=\sum_{v\in\Omega_{K}}\operatorname{inv}_{v}(b(x))$ is the Brauer-Manin pairing of $x$ and $b$. For a finite subgroup $B\subset\operatorname{Br}(U)$, we write $U(\mathbf{A}_{K})^{B}$ for the set of adelic points on $U$ which pair trivially with all elements in $B$. Using this pairing together with character orthogonality, we can give an alternative description of the Tamagawa constant:
@@
In [Pey95], Peyre used a variant of the $\alpha$-constant, which is used in various articles on Manin’s conjecture such as the work on toric varieties by Salberger, Pieropan and Schindler [Sal98, PiSc23]. Rather than taking an exponential integral over the dual of the effective cone, this constant is instead defined as the volume of a slice of this cone. We define a generalization of Peyre’s $\alpha$-constant, which we will use in [Moe25toric] to prove Conjecture 1.4 for toric pairs.
@@
then a general result on cones [BaTs95, Proposition 2.4.4] implies
@@ 9.1. Compatibility with conjecture for Campana points
First we show that Conjecture 1.4 agrees with the conjecture [PSTVA21, Conjecture 1.1] by Pieropan, Smeets, Tanimoto and Várilly-Alvarado with the modified prediction for the leading constant given by Chow, Loughran, Takloo-Bighash and Tanimoto [CLTT24, Conjecture 8.3].
Let $(X,D_{\mathbf{m}})$ be a Campana pair where $X$ is a smooth projective Fano variety over a number field such that the support of $D_{\mathbf{m}}$ is a strict normal crossings divisor, and let $(X,M)$ be the smooth split pair corresponding to the Campana points on this Campana pair. Let $S$ be a finite set of places and let $\mathcal{X}$ be a regular integral model of $X$ over $\mathcal{O}_{S}$. Let $(\mathcal{X},\mathcal{M})$ be the induced integral model of $(X,M)$ and let $(\mathcal{X},\mathcal{D}_{\mathbf{m}})$ be the corresponding Campana pair. By Proposition 6.6 and Proposition 6.10 it follows that $a((X,M),L)=a((X,D_{\mathbf{m}}),L)$ and $b(K,(X,M),L)=b(K,(X,D_{\mathbf{m}}),L)$ for any big and nef divisor class $L$, where the right hand sides are defined as in [PSTVA21, Conjecture 1.1]. Therefore Conjecture 1.4 agrees with [PSTVA21, Conjecture 1.1], up to a differing prediction for the leading constant. Assume as in [CLTT24] that the support of $A=a((X,M),L)L+K_{X}+D_{\mathbf{m}}$ is contained in the support of $D_{\mathbf{m}}$. Then $(X^{\circ},M^{\circ})=(X,M^{\circ})$ is the pair corresponding to Darmon points on $(X\setminus\operatorname{Supp}(A),D_{\mathbf{m}}|_{X\setminus\operatorname{Supp}(A)})$, where $\operatorname{Supp}(A)$ is the support of $A$. In particular, it follows that $\operatorname{Br}_{1}(X,M^{\circ})=\operatorname{Br}_{1}((X,D_{\mathbf{m}}),L)$, where the right hand side is defined as in [CLTT24, Definition 8.1]. Furthermore, a direct calculation of the $\alpha$-constants show that $\alpha((X,M),L)=\alpha((X,D_{\mathbf{m}}),L)$, where the right hand side is defined as in [PSTVA21, §3.3]. These equalities together combine imply that our prediction of the leading constant agrees with the one given in [CLTT24, Conjecture 8.3].
@@ 9.2. Darmon points on vector group compactifications
In [Ito25, Theorem 1.2], Ito proves Conjecture 1.4 for Darmon points on vector group compactifications, where the divisors are chosen to be in the complement of the vector group, and $L$ is assumed to be adjoint rigid with respect to the pair. Although the expression for the leading constant given in his theorem does not precisely match with our description of the leading constant, a direct analogue of [PSTVA21, Lemma 9.3] shows that they are the same.
@@ 9.3. Darmon points and generalized Fermat equations
In [Ara25], Arango-Piñeros counted Darmon points of bounded height on projective space, motivated by the study of primitive solutions of generalized Fermat equations
@@
The Campana pair that he studied is $(\mathbb{P}^{1}_{\mathbb{Z}},(1-\tfrac{1}{a})D_{1}+(1-\tfrac{1}{b})D_{2}+(1-\tfrac{1}{c})D_{3}),$ where $D_{1},D_{2},D_{3}$ are the points $(0:1),(1:1),(1:0)$ and the integers $a,b,c$ are as above. The $\mathbb{Z}$-Darmon points on this pair are the primitive tuples $(x:y)$ such that $|x|$ is an $a$-th power, |xy||x-y| is a $b$-th power and $|y|$ is a $c$-th power. In [Ara25, Theorem 1.2], he proves that the number of these points up to Weil height $B$ tends to

### v3

@@ 6. Quasi-Campana points and the log-canonical divisor
In this section we specialize the Conjecture 1.4 to Campana points, weak Campana points and Darmon points as defined in Definition 3.24. In particular, we will clarify the relation of Conjecture 1.4 with the conjecture [37, Conjecture 1.1] on Campana points of bounded height. In order to uniformly discuss these different kinds of points, we introduce the notion of a quasi-Campana pair.
@@
Let $(X,M)$ be a smooth proper pair over a field $K$. Assume that there exists a log-canonical class $K_{(X,M),\log}$ for $(X,M)$. Let $L$ be a big and nef $\mathbb{Q}$-divisor class on $X$. Then
@@
The assumption on the effective cone in the Proposition is satisfied for many varieties $X$, including Mori dream spaces such as Fano varieties [7, Corollary 1.3.2].
@@
This conjecture is a related to a conjecture by Campana [10, Conjecture 9.10] on Campana rational connectedness for Campana pairs (see also [12, Conjecture 1.4]), which implies Conjecture 6.13 for any pair $(X,M)$ corresponding to the Campana points on $(X,D_{\mathbf{m}})$.
@@ 7.1. The fundamental group
We will introduce the fundamental group of a smooth pair using the valuations $v_{(\mathbf{m},c)}$ introduced in Definition 4.12. For a smooth variety $X$ over a field $K$, the étale fundamental group $\pi_{1}(X)$ admits a description using the valuations of the function field. Indeed, if $f\colon Y\rightarrow X$ is a finite morphism of smooth varieties, then purity of the ramification locus [48, Tag 0BMB] implies that $f$ is étale if and only if it is unramified at every prime divisor $D$ of $X$.
@@
If $(X,M)$ is a smooth pair corresponding to Darmon points, then [30, Lemma 8.7.] and purity of the branch locus [48, Tag 0BMB] together imply that the fundamental group $\pi_{1}(X,M)$ is the étale fundamental group of the corresponding root stack.
@@ 7.2. The Brauer group of a pair
In [29], Mitankin, Nakahara and Streeter introduced the Brauer group $\operatorname{Br}(X,D_{\mathbf{m}})$ for Campana pairs $(X,D_{\mathbf{m}})$ in their study of Darmon points. This Brauer group was then used by Chow, Loughran, Takloo-Bighash and Tanimoto in their prediction [13, Conjecture 8.3] for the leading constant in the analogue of Manin’s conjecture for Campana points. This group is defined as
@@
where $\partial_{D_{i}}$ is the Witt residue as defined in [14, Definition 1.4.11(ii)]. Note that this is the Brauer group of the corresponding root stack [29, Remark 3.15].
@@
is $\partial_{\mathbf{m},c}=\gcd(\mathbf{m})\cdot\partial_{\mathbf{m}/\gcd(\mathbf{m}),c}$, where $K_{(\mathbf{m}/\gcd(\mathbf{m}),c)}$ is the residue field of the valuation $v_{(\mathbf{m}/\gcd(\mathbf{m}),c)}$ and $\partial_{\mathbf{m}/\gcd(\mathbf{m}),c}$ is the residue corresponding to the valuation $v_{\mathbf{m}/\gcd(\mathbf{m}),c}$, as defined in [14, Definition 1.4.11(ii)].
@@ 7.3. The algebraic Brauer group for pairs over toric varieties
In this section we describe the algebraic Brauer group for pairs over toric varieties in terms of automorphic characters on the open torus, which we will use in Section 9.5 and in [32].
Let $(X,M)$ be a smooth pair such that $X$ is a toric variety over a number field and such that the divisors $D_{1},\dots,D_{n}$ are the torus-invariant prime divisors, defined over the splitting field $E/K$ of the dense torus. Note that such a pair is always divisorial, as any nonempty intersection of torus-invariant prime divisors is a toric variety itself and thus irreducible. By using the correspondence between Brauer group elements on a torus and automorphic characters as in [28], we will describe the algebraic Brauer group $\operatorname{Br}_{1}(X,M)$ in terms of characters if $X$ is a (rational) toric variety and the divisors are the torus-invariant divisors.
@@
In particular, given an automorphic character $\chi\colon U(\mathbf{A}_{K})\rightarrow S^{1}$, we obtain an automorphic character $\chi_{\mathbf{m}}\colon T_{\mathbf{m}}(\mathbf{A}_{K})\rightarrow S^{1}$ for every $\mathbf{m}\in\mathbb{N}^{n}_{\textnormal{red}}/G$ (here we recall that $\mathbb{N}^{n}_{\textnormal{red}}$ is the set of all $\mathbf{m}$ such that $\mathcal{C}_{\mathbf{m}}\neq\emptyset$). The following lemma generalizes [45, Lemma 3.25].
@@
Here $\B(U)=\ker(\operatorname{Br}_{1}(U)\rightarrow\prod_{v\in\Omega_{K}}\operatorname{Br}_{1}(U_{K_{v}}))$, which is trivial when $X$ is rational. This isomorphism identifies the Brauer–Manin pairing with the pairing between $U(\mathbf{A}_{K})$ and the corresponding automorphic characters.
@@
For a torus $T$ over $K$ we write $\operatorname{Br}_{e}U=\{b\in\operatorname{Br}_{1}U\mid 1_{T}^{*}b=0\}$, where $1_{T}\colon\operatorname{Spec}K\rightarrow T$ is the identity element, so that we have $\operatorname{Br}_{1}T\cong\operatorname{Br}K\times\operatorname{Br}_{e}T$. We similarly write $\operatorname{Br}_{e}(X,M)=\operatorname{Br}(X,M)\cap\operatorname{Br}_{1}U$.
@@
These observations combine with [28, Lemma 4.7] to give a commutative diagram
@@
Finally, the triviality of $\B(U)$ when $X$ is rational is proven in [28, Corollary 4.6]. ∎
@@ 8. The leading constant
In this section we give a description for the leading constant in Conjecture 1.4 when the divisor class $L$ is adjoint rigid and the fundamental group $\pi(X_{\overline{K}},M)$ of the pair is abelian. Our expression for the leading constant will involve the Fujita invariant and the $b$-invariant, as well as analogues of the $\alpha$-constant and the Tamagawa constant for pairs. In Manin’s conjecture, the complement of the support of the adjoint divisor is used to define several of these invariants.
The $\alpha$-constant and the Tamagawa constant in Manin’s conjecture, as formulated in [6], are defined using the complement $X^{\circ}$ of the adjoint divisor. Our conjecture replaces the role of this open subvariety with a pair $(X^{\circ},M^{\circ})\subset(X,M)$.
@@
where $\tau_{X,v}$ is the measure on $X(K_{v})$ as defined in [11, §2.1.8] induced by the chosen metrisation of $K_{X}$. Note that the measure $\tau_{(X^{\circ},M^{\circ})}$ can be viewed as an analogue of the Tamagawa measure defined in [11, Theorem 1.1].
As in [13, Conjecture 8.3], we define the Tamagawa constant as a sum over Brauer classes in $\operatorname{Br}(X^{\circ},M^{\circ})/\operatorname{Br}K$.
@@
Similarly to the constant in [13, Conjecture 8.3], it is not clear whether the sum defining the Tamagawa constant converges, and if it does, whether it converges absolutely. In the case that the sum does not converge absolutely, the sum $\sum_{b\in\operatorname{Br}(X^{\circ},M^{\circ})/\operatorname{Br}K}\hat{\tau}(b)$ should be interpreted as the limit
@@
where $\operatorname{ev}_{b}(x)=\sum_{v\in\Omega_{K}}\operatorname{inv}_{v}(b(x))$ is the Brauer–Manin pairing of $x$ and $b$. For a finite subgroup $B\subset\operatorname{Br}(U)$, we write $U(\mathbf{A}_{K})^{B}$ for the set of adelic points on $U$ which pair trivially with all elements in $B$. Using this pairing together with character orthogonality, we can give an alternative description of the Tamagawa constant:
@@
In [34], Peyre used a variant of the $\alpha$-constant, which is used in various articles on Manin’s conjecture such as the work on toric varieties by Salberger, Pieropan and Schindler [41, 35]. Rather than taking an exponential integral over the dual of the effective cone, this constant is instead defined as the volume of a slice of this cone. We define a generalization of Peyre’s $\alpha$-constant, which we use in [32] to prove Conjecture 1.4 for toric pairs.
@@
then a general result on cones [4, Proposition 2.4.4] implies
@@ 9.1. Compatibility with conjecture for Campana points
First we show that Conjecture 1.4 agrees with the conjecture [37, Conjecture 1.1] by Pieropan, Smeets, Tanimoto and Várilly-Alvarado with the modified prediction for the leading constant given by Chow, Loughran, Takloo-Bighash and Tanimoto [13, Conjecture 8.3].
Let $(X,D_{\mathbf{m}})$ be a Campana pair where $X$ is a smooth projective Fano variety over a number field such that the support of $D_{\mathbf{m}}$ is a strict normal crossings divisor, and let $(X,M)$ be the smooth split pair corresponding to the Campana points on this Campana pair. Let $S$ be a finite set of places and let $\mathcal{X}$ be a regular integral model of $X$ over $\mathcal{O}_{S}$. Let $(\mathcal{X},\mathcal{M})$ be the induced integral model of $(X,M)$ and let $(\mathcal{X},\mathcal{D}_{\mathbf{m}})$ be the corresponding Campana pair. By Proposition 6.6 and Proposition 6.10 it follows that $a((X,M),L)=a((X,D_{\mathbf{m}}),L)$ and $b(K,(X,M),L)=b(K,(X,D_{\mathbf{m}}),L)$ for any big and nef divisor class $L$, where the right hand sides are defined as in [37, Conjecture 1.1]. Therefore Conjecture 1.4 agrees with [37, Conjecture 1.1], up to a differing prediction for the leading constant. Assume as in [13] that the support of $A=a((X,M),L)L+K_{X}+D_{\mathbf{m}}$ is contained in the support of $D_{\mathbf{m}}$. Then $(X^{\circ},M^{\circ})=(X,M^{\circ})$ is the pair corresponding to Darmon points on $(X\setminus\operatorname{Supp}(A),D_{\mathbf{m}}|_{X\setminus\operatorname{Supp}(A)})$, where $\operatorname{Supp}(A)$ is the support of $A$. In particular, it follows that $\operatorname{Br}_{1}(X,M^{\circ})=\operatorname{Br}_{1}((X,D_{\mathbf{m}}),L)$, where the right hand side is defined as in [13, Definition 8.1]. Furthermore, a direct calculation of the $\alpha$-constants show that $\alpha((X,M),L)=\alpha((X,D_{\mathbf{m}}),L)$, where the right hand side is defined as in [37, §3.3]. These equalities together combine imply that our prediction of the leading constant agrees with the one given in [13, Conjecture 8.3].
@@ 9.2. Darmon points on vector group compactifications
In [25, Theorem 1.2], Ito proves Conjecture 1.4 for Darmon points on vector group compactifications, where the divisors are chosen to be in the complement of the vector group, and $L$ is assumed to be adjoint rigid with respect to the pair. Although the expression for the leading constant given in his theorem does not precisely match with our description of the leading constant, a direct analogue of [37, Lemma 9.3] shows that they are the same.
@@ 9.3. Darmon points and generalized Fermat equations
In [1], Arango-Piñeros counted Darmon points of bounded height on projective space, motivated by the study of primitive solutions of generalized Fermat equations
@@
The Campana pair that he studied is $(\mathbb{P}^{1}_{\mathbb{Z}},(1-\tfrac{1}{a})D_{1}+(1-\tfrac{1}{b})D_{2}+(1-\tfrac{1}{c})D_{3}),$ where $D_{1},D_{2},D_{3}$ are the points $(0:1),(1:1),(1:0)$ and the integers $a,b,c$ are as above. The $\mathbb{Z}$-Darmon points on this pair are the primitive tuples $(x:y)$ such that $|x|$ is an $a$-th power, |xy||x-y| is a $b$-th power and $|y|$ is a $c$-th power. By [1, Theorem 1.2], the number of these points up to Weil height $B$ tends to